OpenStudy (anonymous):

Explain how you would solve the problem below: Nathan has $16 more than Jeff. Jeff has $10 more than Brad. Together the Boys have $66. How much money does Brad have?

OpenStudy (lgbasallote):

let x = Nathan's money y = Jeff's money z = Brad's money x = 16 + y y = 10 + z x+ y + z = 66 Substitute x = 16 + y to the 3rd equation... 16 + y + y + z = 66 16 + 2y + z = 66 2y + z = 50 Use systems of linear equations... 2y + z =50 y - z = 10 3y = 60 y = $20 z = $10 x = $36

OpenStudy (anonymous):

Thank you